OpenStudy (lichking):

Medal to the best answer. If, 2+3=8, 3+7=27, 4+5=32, 5+8=60, 6+7=72, 7+8=??

OpenStudy (anonymous):

umm

OpenStudy (lichking):

yes?

OpenStudy (anonymous):

2*4=8 3*9 = 27 4*8 = 32 5*12=60 ..... now see the 2nd term 4 ,3+1 =4 2nd is 9 ,7+2 =9 3rd is 8,5+3=8 4th 12,8+4 =12 so its adding 1 more number n multiplying by first one sooo 7+8 is 6th term 8+6 =14 14*7 = 98

OpenStudy (lichking):

correct